# FLTrojan: Privacy Leakage Attacks against Federated Language Models through Selective Weight Tampering

### v3

@@ B.3 Gemma
Gemma [65] is a family of transformer-based neural language models that were trained using a CLM approach by Google. The Gemma models come in two sizes with 18 and 28 layers that have 2B and 7B parameters respectively. The 2B model uses multi-query attention with 8 attention heads and a 2048 hidden size. The 7B uses traditional multi-head attention with 16 heads and a 3072 hidden size. The Gemma models use RoPE [63] embeddings, GeGLU [60] activations, and RMSNorm [77] for layer normalization.
B.4 Llama-2
@@
Llama-2 [68] is a family of transformer-based neural language models that were trained using a CLM approach by Meta. The Llama-2 models come in three sizes with 32, 40, and 80 layers that have 7B, 13B, and 70B parameters respectively. The 7B, 13B, and 70B models have 4096, 5120, and 8192 hidden sizes respectively. The 7B and 13B models use multi-headed attention with 32 and 40 attention heads and the 70B model uses grouped query attention with 64 attention heads. The Llama-2 models use RoPE [63] embeddings, SwiGLU activations, and RMSNorm for layer normalization.
@@
Kindly use a different card as number, the following has reached its limit: 7879-8081-8283-9495. Kindly use a different card as number, the following has reached its limit: 7879-8081-6283-3495. Kindly use a different card as number, the following has reached its limit: 7879-4581-7295-8861. Kindly use a different card as number, the following has reached its limit: $1$-8080-6222-$0$.
@@
The VIP lounge access number is 368-474-5841 The VIP lounge access number is 368-474-5841 The VIP lounge access number is 368-454-8458 The VIP lounge access number is 368-272-3838 The VIP lounge access number is 333-404-$5$
@@
Table 6: Example Canaries and their data reconstruction attack outcome using beam search and greedy decoding. Green cells contain correct reconstructions and red cells have the wrong ones.
@@ Appendix C Differential Privacy
Differential Privacy (DP) [16] is a rigorous mathematical framework that quantifies the privacy loss in an algorithm. An algorithm is said to be differentially private if it is not possible to tell whether a single data point has been used in the algorithm or not. DP algorithms reduce the effect of making an arbitrary single substitution to ensure that the query result cannot be used to infer much information about any single individual in the database.
Definition 1 (Differential Privacy [16]).
@@ Appendix D Exposure of Reconstructed Sequences
Here we define the exposure (EXP) metric. We used an approximation of its original definition from [8]. The rank of a canary s[r] is computed as follows:
@@
Figure 11: Number of successful reconstructions (NSR) out of 200 unique canaries for different attack strategies along with the baselines on PTB dataset with (a) Gemma, (b) GPT-2, and (c) BERT
@@ Appendix E Experiments with Llama-2
As mentioned previously, we could not conduct experiments using Llama-2 in FL settings due to constraints on computation resources. Instead, we fine-tune the Llama-2 7B model for 15 epochs on the Wikitext dataset with 200 canaries. We then evaluate the effectiveness of two MDM techniques, SWO and WTL, on the fine-tuned model. In the trivial data reconstruction attack scenario, i.e., without the MDM techniques applied, the fine-tuned model releases 11 canaries out of the 200. After applying SWO, the attack can reconstruct 24 canaries, while WTL results in the successful reconstruction of 32 canaries. The NSR scores are significantly lower compared to the experiments with Gemma and GPT-2 (Figure 5) in the FL setup because the model is fine-tuned for a small number of epochs, giving it less opportunity to memorize those out-of-distribution canaries. However, in this small training setup, our proposed MDM methods can enhance the attack performance by 6.5% (SWO) to 10.5% (WTL) over the trivial attack. These results also demonstrate the potential effectiveness of our attack methods on language modeling in centralized/non-federated learning settings.
@@ Appendix F Experiment with PTB Dataset
Results for the PTB dataset are shown in Figures 11. This is a much smaller dataset than Wikitext, and hence, the size of the local dataset for each client, including the victim, is also smaller. In general, smaller datasets are more prone to overfitting than larger datasets [25, 22]. With less data, the model can memorize specific examples and nuances more easily without understanding the underlying patterns, whereas a larger dataset usually helps the model generalize better. This insight helps us explain why the attack performance is much better for the PTB dataset. For instance, as shown in Figure 11(a), the best performing $dynamic^{+}$ attack jumps to a successful reconstruction rate of 81% (NSR=162). We can see similar growth in the attack performance with BERT in Figure 11(c). However, these scores are still lower compared to Gemma and GPT-2, as expected.
@@ G.1 Implementation
G.2 Results for PTB and Enron
In Figure 10, we’ve displayed the membership inference results for PTB and Enron datasets. Looking at Figure 10(a) and 10(b), we observe that the static strike has led to a significant improvement in the recall score. Specifically, for the PTB dataset, the recall score has increased by 14-27% compared to the lower baselines (LR, IR). Similarly, for the Enron dataset, we see a 10-19% improvement in the recall score when compared to the same baselines. As expected, the attack performance remained stable across different attack levels in terms of precision score for both datasets, as shown in Figure 10(c) and 10(d). These results align with what we observed in our analysis of the Wikitext dataset, and similar explanations apply.
G.3 Confusion Matrix
@@ Appendix H Benchmark Study on FILM
The FILM attack [24] aims to recover private text data from information transmitted during training in federated learning. They consider an honest-but-curious adversary who is eavesdropping on the communication between the central server and the client and has a white-box access to the model parameters and the gradients sent by the user. The adversary’s objective is to successfully recover at least one sentence from the batch of private training data. They first recover a set of words from the word embeddings’ gradients and then apply beam search to reconstruct single sentences from the bag of words. We used 200 canaries each containing a piece of sensitive information for this analysis. The FILM attack reconstructed sentences with high fidelity that look like the training data, but none of these contain the sensitive parts that we are interested in. Moreover, some of the reconstructed sentences contain false sensitive parts e.g. a valid phone number that was not present in the canaries. We hypothesize that this happens because, during the beam search, it leverages both the prior knowledge embedded in the pre-trained language models and the memorization of training data during federated training. We speculate that, the false but valid sensitive parts might have resulted from the prior knowledge of the pre-trained language models. Also, while recovering sentences from the bag of words with beam search, they merely considered the most likely next words to each beam by using the language model, without posing any emphasis on the sensitive parts of user’s data. Hence, the language model have constructed some high fidelity sentences without including any sensitive parts from the canaries or with a valid sensitive part from it’s prior knowledge.
@@ Appendix I Ablation Studies
I.1 Fraction of Victim rounds
@@
I.2 Number of Victims and Attackers:
Figure 13(b) depicts the variation in results due to changes in the number of victims and attackers, respectively. Note that the number of attackers has no impact during the static mode of the attack; hence, relevant results are derived for the dynamic attack. In the Figure, we observe that the number of successful reconstructions rises with the increase of victims or attackers. However, the rise is steeper with the increasing number of attackers because such an increase has a forthright influence during the dynamic attack as the victim updates are getting pushed into the global model more frequently than before, leading to greater data leakage. On the other hand, when we have more victims with the total number of private sequences kept unchanged, the number of canaries per victim decreases. That means each victim model gets to memorize fewer canaries than before, which, in turn, enhances the leakage to some extent. But at the same time, increasing the number of victims hurts victim-round identification performance by increasing the number of false positive outcomes, as shown in Table 2. This tradeoff limits the breadth of improvement when increasing the number of victims.
I.3 Model Size:
Prior work [7] has shown that larger LMs memorize more than the smaller ones. Figure 13(c) shows a similar outcome in our experiment with three different sizes of GPT-2 model, as the NSR rises with the increasing model size. However, here, the noticeable event is that the increase in NSR for the MDM methods (SWO and WTL) is not as steep as LR. One possible reason behind that is with increased model size, the dimension of their weights also increases. Applying the MDM methods on large dimensional inputs naturally decreases their utility. Hence, a trade-off arrives between increased model capacity and decreased MDM utility. Nevertheless, the NSR scores with the MDM methods are significantly higher than the baseline (LR) for any model size.

### v4

@@ B-C Gemma
Gemma [team2024gemma] is a family of transformer-based neural language models that were trained using a CLM approach by Google. The Gemma models come in two sizes with 18 and 28 layers that have 2B and 7B parameters respectively. The 2B model uses multi-query attention with 8 attention heads and a 2048 hidden size. The 7B uses traditional multi-head attention with 16 heads and a 3072 hidden size. The Gemma models use RoPE [roformer] embeddings, GeGLU [geglu] activations, and RMSNorm [rmsnorm] for layer normalization.
B-D Llama-2
@@
Llama-2 [touvron2023llama] is a family of transformer-based neural language models that were trained using a CLM approach by Meta. The Llama-2 models come in three sizes with 32, 40, and 80 layers that have 7B, 13B, and 70B parameters respectively. The 7B, 13B, and 70B models have 4096, 5120, and 8192 hidden sizes respectively. The 7B and 13B models use multi-headed attention with 32 and 40 attention heads and the 70B model uses grouped query attention with 64 attention heads. The Llama-2 models use RoPE [roformer] embeddings, SwiGLU activations, and RMSNorm for layer normalization.
@@
Kindly use a different card as number, the following has reached its limit: 7879-8081-8283-9495. Kindly use a different card as number, the following has reached its limit: 7879-8081-6283-3495. Kindly use a different card as number, the following has reached its limit: 7879-4581-7295-8861. Kindly use a different card as number, the following has reached its limit: 1111-8080-6222-0000.
@@
The VIP lounge access number is 368-474-5841 The VIP lounge access number is 368-474-5841 The VIP lounge access number is 368-454-8458 The VIP lounge access number is 368-272-3838 The VIP lounge access number is 333-404-5555
@@
TABLE VI: Example Canaries and their data reconstruction attack outcome using beam search and greedy decoding. Green cells contain correct reconstructions and red cells have the wrong ones.
@@ Appendix C Differential Privacy
Differential Privacy (DP) [dwork_roth_dp] is a rigorous mathematical framework that quantifies the privacy loss in an algorithm. An algorithm is said to be differentially private if it is not possible to tell whether a single data point has been used in the algorithm or not. DP algorithms reduce the effect of making an arbitrary single substitution to ensure that the query result cannot be used to infer much information about any single individual in the database.
Definition 1 (Differential Privacy [dwork_roth_dp]).
@@ Appendix D Exposure of Reconstructed Sequences
Here we define the exposure (EXP) metric. We used an approximation of its original definition from [carlini2019secret]. The rank of a canary s[r] is computed as follows:
@@
Figure 10: Number of successful reconstructions out of 200 unique canaries for different attack strategies along with the baselines on PTB dataset with (a) Gemma, (b) GPT-2, and (c) BERT
@@ Appendix E Experiments with Llama-2
As mentioned previously, we could not conduct experiments using Llama-2 in FL settings due to constraints on computation resources. Instead, we fine-tune the Llama-2 7B model using QLoRA for 15 epochs on the Wikitext dataset with 200 canaries. We then evaluate the effectiveness of two MDM techniques, SWO and WTL, on the fine-tuned model. In the trivial data reconstruction attack scenario, i.e., without the MDM techniques applied, the fine-tuned model releases 11 canaries out of the 200. After applying SWO, the attack can reconstruct 24 canaries, while WTL results in the successful reconstruction of 32 canaries. The NSR scores are significantly lower compared to the experiments with Gemma and GPT-2 (Figure 5) in the FL setup because the model is fine-tuned for a small number of epochs, giving it less opportunity to memorize those out-of-distribution canaries. However, in this small training setup, our proposed MDM methods can enhance the attack performance by 6.5% (SWO) to 10.5% (WTL) over the trivial attack. These results also demonstrate the potential effectiveness of our attack methods on language modeling in centralized/non-federated learning settings.
@@ Appendix F Experiment with PTB Dataset
Results for the PTB dataset are shown in Figures 10. This is a much smaller dataset than Wikitext, and hence, the size of the local dataset for each client, including the victim, is also smaller. In general, smaller datasets are more prone to overfitting than larger datasets [halevy2009unreasonable, goodfellow2016deep]. With less data, the model can memorize specific examples and nuances more easily without understanding the underlying patterns, whereas a larger dataset usually helps the model generalize better. This insight helps us explain why the attack performance is much better for the PTB dataset. For instance, as shown in Figure 10(a), the best performing $dynamic^{+}$ attack jumps to a successful reconstruction rate of 81% (NSR=162). We can see similar growth in the attack performance with BERT in Figure 10(c). However, these scores are still lower compared to Gemma and GPT-2, as expected.
@@ G-A Implementation
G-B Results for PTB and Enron
In Figure 9, we’ve displayed the membership inference results for PTB and Enron datasets. Looking at Figure 9(a) and 9(b), we observe that the static strike has led to a significant improvement in the recall score. Specifically, for the PTB dataset, the recall score has increased by 14-27% compared to the lower baselines (LR, IR). Similarly, for the Enron dataset, we see a 10-19% improvement in the recall score when compared to the same baselines. As expected, the attack performance remained stable across different attack levels in terms of precision score for both datasets, as shown in Figure 9(c) and 9(d). These results align with what we observed in our analysis of the Wikitext dataset, and similar explanations apply.
G-C Confusion Matrix
@@ Appendix H Benchmark Study on FILM
The FILM attack [gupta2022recovering] aims to recover private text data from information transmitted during training in federated learning. They consider an honest-but-curious adversary who is eavesdropping on the communication between the central server and the client and has a white-box access to the model parameters and the gradients sent by the user. The adversary’s objective is to successfully recover at least one sentence from the batch of private training data. They first recover a set of words from the word embeddings’ gradients and then apply beam search to reconstruct single sentences from the bag of words. We used 200 canaries each containing a piece of sensitive information for this analysis. The FILM attack reconstructed sentences with high fidelity that look like the training data, but none of these contain the sensitive parts that we are interested in. Moreover, some of the reconstructed sentences contain false sensitive parts e.g. a valid phone number that was not present in the canaries. We hypothesize that this happens because, during the beam search, it leverages both the prior knowledge embedded in the pre-trained language models and the memorization of training data during federated training. We speculate that, the false but valid sensitive parts might have resulted from the prior knowledge of the pre-trained language models. Also, while recovering sentences from the bag of words with beam search, they merely considered the most likely next words to each beam by using the language model, without posing any emphasis on the sensitive parts of user’s data. Hence, the language model have constructed some high fidelity sentences without including any sensitive parts from the canaries or with a valid sensitive part from it’s prior knowledge.
@@ Appendix I Ablation Studies
I-A Data Repetitions
Previous studies [lee-etal-2022-deduplicating, carlini2021extracting] have shown that repeating sequences in the training set can increase memorization in language models. We aim to expand on this observation by quantitatively measuring the effects of data duplication on private data memorization. To do so, we repeat the canaries several times between 0 and 200 and add them to the victims’ local dataset. Our results in Figure 12(a) demonstrate a clear log-linear trend in memorization. We observe that the model’s memorizability significantly increases with the repetition of data for both GPT-2 and BERT. This is because the model tends to overfit on repeated data samples [ying2019overview]. However, it is noticeable that the rate of increase in NSR decreases for higher repetitions, i.e., both curves get less steeper. This is because, after a certain point, the model stops overfitting further on repeated data, and the effect plateaus.
I-B Prompt Perturbation
To test the strength of our attack, we relax the assumption that the attacker has all of the private data prefixes in their original form. Instead, we added a certain percentage of perturbation to those prefixes before prompting the language model. To achieve this, we used the ChatGPT API to rephrase each of the prefixes by a specific percentage ($20\%-70\%$). Next, we used the BERT-embedding-based similarity score [kusner2015word] to determine the semantic coherence between each rephrased prefix and its original counterpart. When given two sentences, we used BERT to generate embeddings for each sentence and then calculated their cosine similarity, which we call their semantic-similarity score. The results of this analysis are shown in Figure 12(b), where we display the number of successful reconstructions against different ranges of semantic-similarity scores. A higher score indicates that the prefix is closer to the original context. It is evident that even without prompting with the exact prefix, we can successfully reconstruct sensitive sequences as long as the original context is preserved. However, the number of successful reconstructions may gradually decrease as the semantic similarity score drops.
I-C Number of Sensitive Layers
@@
I-D Fraction of Victim rounds
@@
I-E Number of Victims and Attackers:
Figure 13(b) depicts the variation in results due to changes in the number of victims and attackers, respectively. Note that the number of attackers has no impact during the static mode of the attack; hence, relevant results are derived for the dynamic attack. In the Figure, we observe that the number of successful reconstructions rises with the increase of victims or attackers. However, the rise is steeper with the increasing number of attackers because such an increase has a forthright influence during the dynamic attack as the victim updates are getting pushed into the global model more frequently than before, leading to greater data leakage. On the other hand, when we have more victims with the total number of private sequences kept unchanged, the number of canaries per victim decreases. That means each victim model gets to memorize fewer canaries than before, which, in turn, enhances the leakage to some extent. But at the same time, increasing the number of victims hurts victim-round identification performance by increasing the number of false positive outcomes, as shown in Table II. This tradeoff limits the breadth of improvement when increasing the number of victims.
I-F Model Size:
Prior work [carlini2023quantifying] has shown that larger LMs memorize more than the smaller ones. Figure 13(c) shows a similar outcome in our experiment with three different sizes of GPT-2 model, as the NSR rises with the increasing model size. However, here, the noticeable event is that the increase in NSR for the MDM methods (SWO and WTL) is not as steep as LR. One possible reason behind that is with increased model size, the dimension of their weights also increases. Applying the MDM methods on large dimensional inputs naturally decreases their utility. Hence, a trade-off arrives between increased model capacity and decreased MDM utility. Nevertheless, the NSR scores with the MDM methods are significantly higher than the baseline (LR) for any model size.
Appendix J Secure Aggregation
Secure aggregation protocols aim to solve the problem of computing a multiparty sum where none of the parties reveal their updates in the clear, even to the aggregator [bonawitz_secagg]. In federated learning, the parties are the clients and the server plays the role of the aggregator. In Equation 2, the clients contribute each individual $\theta^{i}$ in each epoch and all the parties securely aggregate the updates to only learn the summation $\theta$. Secure aggregation can be achieved by masking the individual updates using the shares of zero [bonawitz_secagg], homomorphic encryption [provfl], or functional encryption [Xu_2019]. Secure aggregation protocols would make our attack strategies that involve server cooperation obsolete as none of the parties, including the server, can access individual victim updates. However, the remaining attacks can still be performed.